2/5x-1=7/5 how to even start.
-54/3 + 3/2x + 2 ---dont even know where to start.. help please..

Answer by UJIMUSHKU(3)

2/5x-1=7/5

To solve this equation one must isolate X, that means to get X on one side of the equation and everything else on the other side.
1. Add one to both sides of the equation.
2/5x -1 +1 = 7/5 +1
2. Add 7/5 + 1.
2/5x = 12/5
3. Divide both sides by 2/5.
2/5x / 2/5 = 12/5 / 2/5
4. When you divide 2/5x by 2/5 you get x.
5. To divide 12/5 by 2/5 means to multiply by the inverse of the second fraction. 12/5 * 5/2 = 60/10
6. We are left with X = 60/10
now reduce the fraction to get
X = 6

-54/3 + 3/2x + 2
To solve this start by setting the equation equal to zero or
-52/3 + 3/2x +2 = 0. Now you must isolate X to get the answer.
1. Subtract 3/2x from both sides of the equation.
-54/3 +3/2x + 2 -3/2x = -3/2x
2. Add the numbers on the left side of the equation together.
-54/3 + 2/1 = -54/3 + 6/3 = -48/3
Now we have -48/3 = -3/2x
3. Divide both sides of the equation by -3/2 to isolate the x.
-48/3 * -2/3 = -3/2x / -3/2
Next we get 96/9 = X
4. Last reduce the fraction to get your answer.
X = 32/3